OpenStudy (anonymous):

Can anyone anwser this? x2 - 12x = -36

OpenStudy (anonymous):

(x-6)(x-6)

OpenStudy (anonymous):

x2-12x+36=0 factorize x2-6x-6x+36=0 x(x-6)-6(x-6)=0 (x-6) (x-6) =0 x=6 answer

OpenStudy (anonymous):

Thanks!